OpenStudy (anonymous):

A ball is thrown horizontally with a speed of 10 m/s, from the edge of the top of the building of height 19.6m, and later strikes the ground. 1) For how long was the ball in the air? 2)what is the vertical component of the velocity just before the ball strikes the ground? 3) what is the horizontal component of the velocity just before the ball strikes the ground? 4)how far from the base of the building did the ball strike the ground? 5) what is the speed of the ball just before it strikes the ground?

OpenStudy (anonymous):

1) 19.6=.5(9.8)t^2 Solve for t 2) vf^2=2(9.8)(19.6) solve for vf 3) This never changes still 10m/s 4) use your time from part 1 times 10m/s 5) you will need to do Pythagorean theorem with 10m/s and vf

OpenStudy (anonymous):

can you type out the equations you used for these parts and can you clarify #5 please?

OpenStudy (anonymous):

part 4 is just your answer in part 2 times 10 Part 5 is basically saying you have two velocities, one in the x direction and one in the y direction. They make a right angle with each other (vector addition). So, you need to find the long side, do this with the theorem \[\sqrt{side1^2 + side2^2}\]

OpenStudy (anonymous):

thanks for the help, sorry for the late reply i had to run back to my apartment really fast.

OpenStudy (anonymous):

haha no problem